Find the equation of the line in standard form through the points (-3,9) (5,7)
Answer by Vijay Ramrakhiani(2)

To find the slope the formula is m=Y2-Y1 / X2-X1
So here X1 is= -3 and X2 is= 5
And Y1 is= 9 and Y2 is=7
And put the value now in formula the slope will come.
m= 7-9/5-(-3)
m=-3/8
so the slope is -3/8
Now we have to find the Equation:
So the formula is find the equation:
y-Y1=m(X-X1)
put the values
y-9=-3/8(X-(-3))
y-9=-3x/8 - 9/8
y=-3x/8 - 9/8 +9
y=-3x/8 - (9/8-9/1)
y=-3x/8+63/8 this is the equation